# Cohomology of Lie algebroids over Algebraic spaces

### v6

@@ 3.2. Chevalley-Eilenberg-de Rham complex and Koszul-Rinehart resolution
For a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ with a representation $(\mathcal{E},\nabla)$, consider the cochain complex (consists with $\mathcal{O}_{X}$-modules with a $\mathbb{K}_{X}$-linear degree $1$ map), a generalization of the well known Chevalley-Eilenberg-de Rham complex [Mac05]
@@
It is the sheafification of the presheaf of cochain complexes [Wei96]
@@
where $D_{1},\dots,D_{k+1}\in\mathcal{L}(U)$ and $\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U))$, and $\nabla^{U}:=\nabla(U)$ (see 2,3). Notice that, the differential $d$ (i.e. $d^{2}=0)$ is a $\mathbb{K}_{X}$-linear map, but not an $\mathcal{O}_{X}$-module homomorphism $(d(f~{}\omega)=df\wedge\omega+f~{}d\omega$, for $f\in\mathcal{O}_{X}(U),\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U)))$.
@@
which implies $\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E})\cong\Omega^{\bullet}_{\mathcal{L}}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ where $\Omega^{\bullet}_{\mathcal{L}}:=\Omega^{\bullet}_{\mathcal{L}}(\mathcal{O}_{X})$. This complex is equivalent to the standard complex associated with a smooth Lie algebroid [Mac05] or a holomorphic Lie algebroid [Pym13]. As a special case, we get the Chevalley-Eilenberg complex when $\mathcal{L}$ is a Lie algebroid over a point; and when we consider $\mathcal{L}=\mathcal{T}_{X}$ and $\mathcal{E}=\mathcal{O}_{X}$ where $X$ is non-singular (smooth), to recover the de Rham complex $\Omega^{\bullet}_{X}$.
@@
it is the sheafification of the presheaf of cochain complexes [Wei96]
@@
where $\tilde{D}\bar{D_{i}}$ define by the canonical map $\iota_{U}:\mathcal{L}(U)\rightarrow\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$, for all $\tilde{D}\in\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ and $D_{1},\dots,D_{k}\in\mathcal{L}(U)$. Notice that $\partial$ is $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-linear.
When $\mathcal{L}$ is locally free as $\mathcal{O}_{X}$-module, it provides a locally free resolution of $\mathcal{O}_{X}$ (as a left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module), known as the Koszul-Rinehart resolution. That is, it is given by the (augmented) chain complex of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules
@@
In the next section we consider analogous resolutions for $\mathcal{O}_{X}$ in the category of left $\mathcal{S}_{{O}_{X}}\mathcal{L}$-comodules, where we replace $\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$ by dual Koszul-Rinehart complex $\widetilde{\mathscr{K}}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$ (sheafifying the local counterpart as described in [KP11] accordingly).
@@
In [Pym13, BMRT15, Bru17], the Lie algebroid cohomology have considered for a special kind of Lie algebroid $\mathcal{L}$ (consists by coherent sheaf or locally free sheaf of finite rank) over a Noetherian separated scheme or a complex manifold $(X,\mathcal{O}_{X})$. The Lie algebroid cohomology of $\mathcal{L}$ with values in a $\mathcal{L}$-module $\mathcal{E}$ (a coherent $\mathcal{O}_{X}$-module) is the hypercohomology of the complex $(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}\otimes_{\mathcal{O}_{X}}\mathcal{E},$ $d)$. This appears as a special case (5) of the complex described in (4).
@@
We now describe Lie algebroid cohomology as a derived functor.
@@
We have a canonical morphism between these two presheaves of cochain complexes as $\mathbb{K}_{X}$-modules, which is a local isomorphism or stalk-wise isomorphism, using results from [Rin63] as local descriptions (applying on each special open neighborhoods $U_{x}$’s of $x\in X$). Thus, the associated sheaves of cochain complexes are isomorphic as $\mathbb{K}_{X}$-modules (using Lemma $3.2$ and Lemma $3.3$ from [MS22]), i.e.
@@
Note 3.6.
One can find the proof for a locally free Lie algebroid $\mathcal{L}$ of finite rank over a Noetherian separated scheme $(X,\mathcal{O}_{X})$ discussed in [Bru17], by using the ideas of [Rin63] on the level of stalks $\mathcal{L}_{x}$ (since $\mathcal{L}_{x}$ is a $(\mathbb{K},\mathcal{O}_{X,x})$-Lie-Rinehart algebra, projective (in fact, free) module over $\mathcal{O}_{X,x}$ for all $x\in X)$.
@@
Consider the path algebroid $\mathcal{P}_{X}$ for a non-singular special $a$-space $(X,\mathcal{O}_{X})$ and its universal enveloping algebroid $\mathbb{D}_{X}$ (see Remark 2.12). Note that a module over the free Lie algebroid $\mathcal{P}_{X}$ can be viewed as a vector bundle with connection and a $\mathbb{D}_{X}$-module describe some system of PDE on the space of paths (see [Kap07] for details).
@@
Let $\mathcal{A}$ be a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra [MS22] of finite or infinite type (for examples consider Note 2.16). Then the Lie algebroid $\mathscr{P}(\mathcal{A})$ (sheaf of primitive elements, see Note 2.15) is a locally free $\mathcal{O}_{X}$-module (of finite or infinite rank accordingly). Thus, we have $\mathcal{A}\cong\mathscr{U}(\mathcal{O}_{X},\mathscr{P}(\mathcal{A}))$ [MS22]. Therefore, applying Theorem 3.5 for $\mathcal{L}=\mathscr{P}(\mathcal{A})$ provides the isomorphism
@@
In complex geometry (algebraic geometry), for a cochain complex of coherent sheaves $\mathcal{F}^{\bullet}$ over an analytic space (algebraic variety), we compute the hypercohomology $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$ via Čech cohomology $\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$ associated with some good open cover $\mathcal{U}=\{U_{i}\}_{i}$ of $X$ (since $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})\cong\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$, given by a canonical isomorphism). Specifically, we provide a good open cover by connected Stein spaces (affine varieties) and we use Leray’s theorem [Dan89, Ram05] and Cartan-Serre’s vanishing theorem [For11] for sheaf cohomology.
@@
In both cases we have $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong\check{H}^{\bullet}(\mathcal{U},\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E}))$. Since over Stein spaces (affine varieties) $U_{i}$, $\mathcal{L}(U_{i})$ is a projective $\mathcal{O}_{X}(U_{i})$-module [Mor13, Wei13], thus
@@
In [BMRT15], the Lie algebroid (hyper)cohomology of holomorphic (algebraic) Lie algebroids over a complex manifold (smooth scheme) has expressed by Čech cohomology using a good open cover consisting with connected Stein manifolds (affine schemes). But the case associated with analytic spaces is not considered. Also, note that the whole ideas works for the classical case of Lie algebroids over smooth manifolds (i.e. for smooth Lie algebroids).
More generally, for a cochain complex of quasicoherent sheaves $\mathcal{F}^{\bullet}$ over a Noetherian separated scheme, by considering an affine open cover $\mathcal{U}=\{U_{i}\}_{i}$ of $X$, we get a canonical isomorphism $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})\cong\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$ [Har77]. Thus, we get a similar result for a locally free Lie algebroid over a Noetherian separated scheme.
@@ 3.3. Chern classes:
Lie algebroid (or Lie-Rinehart algebra) connections, curvatures and the associated Chern classes (one of the Characteristic classes) has been studied in the context of differential geometry [Fer02] and algebric geometry [Maa05, Maa07, Maa15]. We review and extend some of the facts in our context as follows.
@@
where $Ext^{1}(L,E,\nabla)$ is the set of equivalence classes of extensions of $L$ by the flat connection $(E,\nabla)$ [Maa15]. This result can be viewed in the algebraic geometry set up (over the affine scheme $SpecR$). Analogously, for a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with a $\mathcal{L}$-module $(\mathcal{E},\nabla)$ we get the isomorphism of $\mathbb{K}$-vector spaces
@@
Note that the Chern class of a finitely generated projective $R$-module $E$ with a $L$-connection $\nabla$ on $E$, defined by the trace of the curvature $R_{\nabla}$ associated with the connection $\nabla$ [Maa07, Maa15]. If $R$ is a regular $\mathbb{K}$-algebra (implies $L:=Der_{\mathbb{K}}(R)$ is a finitely generated projective $R$-module) then by considering the covariant connection on $E$ we get the first Chern class given by trace of the curvature
@@
Analogous result holds for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank over the affine scheme $X=SpecR$. We compute analogous result in the general set up by sheafifying the local descriptions given in [Maa05, Maa07].
@@
The restriction of a $\mathcal{T}_{X}$-connection to the Lie algebroid $\mathcal{L}:=\mathcal{T}_{X}(-logY)$ ($\Omega_{\mathcal{L}}^{1}:=\Omega_{X}^{1}(logY)$) induces a logarithmic connection which is a meromorphic connection with poles along the divisor $Y$. Moreover, if $Y$ is a free divisor, then for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank with a $\mathcal{L}$-connection we get the first Chern class $c_{1}(\mathcal{E})\in\mathbb{H}^{2}(\mathcal{L},\mathcal{O}_{X})$. Note that a $\mathcal{L}$-connection might not exist in general. We can extend the notion of Chern classes for $\mathcal{T}_{X}$ or $\mathcal{T}_{X}(-logY)$ to arbitary Lie algebroids.

### v7

@@ 3.2. Chevalley-Eilenberg-de Rham complex and Koszul-Rinehart resolution
For a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ with a representation $(\mathcal{E},\nabla)$, consider the cochain complex (consists with $\mathcal{O}_{X}$-modules with a $\mathbb{K}_{X}$-linear degree $1$ map), a generalization of the well known Chevalley-Eilenberg-de Rham complex [21]
@@
It is the sheafification of the presheaf of cochain complexes [34, 39]
@@
where $D_{1},\dots,D_{k+1}\in\mathcal{L}(U)$ and $\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U))$, and $\nabla^{U}:=\nabla(U)$ (see the notion in (2), (3)). Notice that, the differential $d$ (i.e. $d^{2}=0)$ is a $\mathbb{K}_{X}$-linear map, but not an $\mathcal{O}_{X}$-module homomorphism ((((satisfies the Leibniz rule: $d(f~{}\omega)=df\wedge\omega+f~{}d\omega$, for $f\in\mathcal{O}_{X}(U),\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U)))$.
@@
which implies $\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E})\cong\Omega^{\bullet}_{\mathcal{L}}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ where $\Omega^{\bullet}_{\mathcal{L}}:=\Omega^{\bullet}_{\mathcal{L}}(\mathcal{O}_{X})$. This complex is equivalent to the standard complex associated with a smooth Lie algebroid [21] or a holomorphic Lie algebroid [32]. In particular, we get the Chevalley-Eilenberg complex when $\mathcal{L}$ is a Lie algebroid over a point; and the de Rham complex $\Omega^{\bullet}_{X}$ when $\mathcal{L}=\mathcal{T}_{X}$ and $\mathcal{E}=\mathcal{O}_{X}$ over a non-singular (or smooth) $a$-space $X$.
@@
it is the sheafification of the presheaf of cochain complexes [34, 39]
@@
where $\tilde{D}\bar{D_{i}}$ define by the canonical map $\iota_{U}:\mathcal{L}(U)\rightarrow\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$, for all $\tilde{D}\in\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ and $D_{1},\dots,D_{k}\in\mathcal{L}(U)$. Notice that the boundary (or chain) map $\partial$ is $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-linear.
When $\mathcal{L}$ is locally free as $\mathcal{O}_{X}$-module, it provides a locally free resolution of $\mathcal{O}_{X}$ as a left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module (implies the stalk-wise exactness), known as the Koszul-Rinehart resolution. That is, it is given by the (augmented) chain complex of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules
@@
Note 3.4.
In the last section we consider analogous resolutions for $\mathcal{O}_{X}$ in the category of left $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$-comodules, where we replace the complex $\mathscr{K}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$ with its dual complex, the dual Koszul-Rinehart complex $\widetilde{\mathscr{K}}_{\mathcal{O}_{X}}^{\bullet}\mathcal{L}$ ((((sheafifying the local counterpart as described in [19] accordingly)))).
@@
In [32, 1, 3], the Lie algebroid cohomology have considered for a special kind of Lie algebroid $\mathcal{L}$ (consists by coherent sheaf or locally free sheaf of finite rank) over a Noetherian separated scheme or a complex manifold $(X,\mathcal{O}_{X})$. The Lie algebroid cohomology of $\mathcal{L}$ with values in a $\mathcal{L}$-module $\mathcal{E}$ (a coherent $\mathcal{O}_{X}$-module) is the hypercohomology of the complex $(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}\otimes_{\mathcal{O}_{X}}\mathcal{E},$ $d)$. This appears as a special case (5) of the complex described in (4).
@@
We now describe Lie algebroid cohomology in terms of a derived functor.
@@
We have a canonical morphism between these two presheaves of cochain complexes as $\mathbb{K}_{X}$-modules, which is a local isomorphism or stalk-wise isomorphism, using results from [34] as local descriptions (applying on each special open neighborhoods $U_{x}$’s of $x\in X$). Thus, the associated sheaves of cochain complexes are isomorphic as $\mathbb{K}_{X}$-modules (using Lemma $3.2$ and Lemma $3.3$ from [29]), i.e.
@@
Note 3.7.
One can find the proof for a locally free Lie algebroid $\mathcal{L}$ of finite rank over a Noetherian separated scheme $(X,\mathcal{O}_{X})$ discussed in [3], by using the ideas of [34] on the level of stalks $\mathcal{L}_{x}$ (since $\mathcal{L}_{x}$ is a $(\mathbb{K},\mathcal{O}_{X,x})$-Lie-Rinehart algebra, projective (in fact, free) module over $\mathcal{O}_{X,x}$ for all $x\in X)$.
@@
Consider the path algebroid $\mathcal{P}_{X}$ for a non-singular special $a$-space $(X,\mathcal{O}_{X})$ and its universal enveloping algebroid $\mathbb{D}_{X}$ (see Remark 2.12). Note that a module over the free Lie algebroid $\mathcal{P}_{X}$ can be viewed as a vector bundle with connection and a $\mathbb{D}_{X}$-module describe some system of PDE on the space of paths (see [18] for details).
@@
Let $\mathcal{A}$ be a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra [29] of finite or infinite type (for examples consider Note 2.16). Then the Lie algebroid $\mathscr{P}(\mathcal{A})$ (sheaf of primitive elements, see Note 2.15) is a locally free $\mathcal{O}_{X}$-module (of finite or infinite rank accordingly). Thus, we have $\mathcal{A}\cong\mathscr{U}(\mathcal{O}_{X},\mathscr{P}(\mathcal{A}))$ [29]. Therefore, applying Theorem 3.6 for $\mathcal{L}=\mathscr{P}(\mathcal{A})$ provides the isomorphism
@@
In complex geometry ((((algebraic geometry)))), for a cochain complex of coherent sheaves $\mathcal{F}^{\bullet}$ over an analytic space ((((algebraic variety)))), we compute the hypercohomology $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$ via Čech cohomology $\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$ associated with some good open cover $\mathcal{U}=\{U_{i}\}_{i}$ of $X$ ((((since $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})\cong\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$, given by a canonical isomorphism). Specifically, we provide a good open cover by connected Stein spaces (affine varieties) and we use Leray’s theorem [8, 33] and Cartan-Serre’s vanishing theorem [11] for sheaf cohomology.
@@
In both cases we have $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong\check{H}^{\bullet}(\mathcal{U},\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E}))$. Since over Stein spaces ((((affine varieties)))) $U_{i}$, $\mathcal{L}(U_{i})$ is a projective $\mathcal{O}_{X}(U_{i})$-module [28, 38], thus we get an isomorphism
@@
Also, note that the whole ideas works for the classical case of Lie algebroids over smooth manifolds (i.e. for smooth Lie algebroids). In [1], the Lie algebroid ((((hyper))))cohomology of holomorphic ((((algebraic)))) Lie algebroids over a complex manifold ((((smooth scheme)))) has expressed by Čech cohomology using a good open cover consisting with connected Stein manifolds ((((affine schemes)))). However, the case associated with analytic spaces is not considered. Additionally, this result can be extended to quasicoherent $\mathcal{O}_{X}$-modules.
More generally, for a cochain complex of quasicoherent sheaves $\mathcal{F}^{\bullet}$ over a Noetherian separated scheme, by considering an affine open cover $\mathcal{U}=\{U_{i}\}_{i}$ of $X$, we get a canonical isomorphism $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})\cong\check{H}^{\bullet}(\mathcal{U},\mathcal{F}^{\bullet})$ [14]. Thus, we get a similar result for a locally free Lie algebroid over a Noetherian separated scheme.
@@ 3.3. Chern classes:
Lie algebroid (or Lie-Rinehart algebra) connections, curvatures and the associated Chern classes (one of the Characteristic classes) has been studied in the context of differential geometry [10] and algebric geometry [23, 24, 25]. We review and extend some of the facts in our context as follows.
@@
where $Ext^{1}(L,E,\nabla)$ is the set of equivalence classes of extensions of $L$ by the flat connection $(E,\nabla)$ [25]. This result can be viewed in the algebraic geometry set up (over the affine scheme $SpecR$). Analogously, for a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with a $\mathcal{L}$-module $(\mathcal{E},\nabla)$ we get the isomorphism of $\mathbb{K}$-vector spaces
@@
Note that the Chern class of a finitely generated projective $R$-module $E$ with a $L$-connection $\nabla$ on $E$, defined by the trace of the curvature $R_{\nabla}$ associated with the connection $\nabla$ [24, 25]. If $R$ is a regular $\mathbb{K}$-algebra (implies $L:=Der_{\mathbb{K}}(R)$ is a finitely generated projective $R$-module) then by considering the covariant connection on $E$ we get the first Chern class given by trace of the curvature
@@
Analogous result holds for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank over the affine scheme $X=SpecR$. We compute analogous result in the general set up by sheafifying the local descriptions given in [23, 24].
@@
The restriction of a $\mathcal{T}_{X}$-connection to the Lie algebroid $\mathcal{L}:=\mathcal{T}_{X}(-logY)$ ((((i.e. $\Omega_{\mathcal{L}}^{1}:=\Omega_{X}^{1}(logY)$) induces a logarithmic connection which is a meromorphic connection with poles along the divisor $Y$. Moreover, if $Y$ is a free divisor, then for a locally free $\mathcal{O}_{X}$-module $\mathcal{E}$ of finite rank with a $\mathcal{L}$-connection we get the first Chern class $c_{1}(\mathcal{E})\in\mathbb{H}^{2}(\mathcal{L},\mathcal{O}_{X})$. Note that a $\mathcal{L}$-connection might not exist in general. We can extend the notion of Chern classes for $\mathcal{T}_{X}$ or $\mathcal{T}_{X}(-logY)$ to arbitary Lie algebroids.